# EXFormer: A Multi-Scale Trend-Aware Transformer with Dynamic Variable Selection for Foreign Exchange Returns Prediction

### v1

@@ 1 Introduction
Exchange rates lie at the core of global finance, shaping trade, capital flows, monetary policy transmission, portfolio allocation, and corporate risk management [fang202430]. Accurate day-ahead forecasts support hedging by multinational firms, guide central bank interventions, and inform systematic trading strategies used by hedge funds [rime2010exchange]. Yet despite decades of research, forecasting daily exchange rate returns remains one of the most persistent challenges in international finance [faust2003exchange, rossi2013exchange]. meese1983empirical demonstrated that structural and time-series models routinely fail to outperform a simple random walk in out-of-sample tests. The Meese–Rogoff puzzle reflects two fundamental difficulties. First, daily returns combine high-frequency noise—microstructure effects, news-driven volatility, liquidity shocks—with slower patterns such as medium-term cycles from interest differentials and carry-trade activity [hasanov2024structural, mueller2017exchange, narayan2022understanding, mao2024unveiling]. Second, it is necessary to model the influence and cross-market interactions of diverse exogenous drivers, such as other currency pairs, stock and commodity indices, and scheduled macroeconomic announcements on the daily returns behavior [stavrakeva2024fundamental]. Traditional econometric models, especially those designed for lower-frequency data, often produce unstable daily forecasts and are prone to overfitting [darvas2025forecasting, iyke2022exchange]. Portfolio-balance theory further highlights the role of cross-asset capital flows and time-varying risk preferences, which conventional models typically overlook [engel2019uncovered]. The core research gap is thus how to jointly model multi-scale temporal dependencies and multi-factor cross-market linkages. Although uncovered interest parity (UIP) implies unpredictability of exchange rate changes, its persistent empirical rejection indicates the presence of time-varying risk premia that more advanced forecasting models may exploit [fang202430, ayitey2023forex].
@@
Recent work increasingly applies machine learning and deep learning to forecasting daily exchange rate returns [plakandaras2015forecasting, sun2020new, neghab2025explaining]. Classical methods such as SVMs and random forests offer limited gains, as they ignore temporal structure [sadeghi2021combined, gyamerah2020long]. RNN-based models including LSTM and GRU networks, better capture sequential dependencies [zhou2025forecasting, cao2020deep] and have been widely tested in financial forecasting (fischer2018deep; gu2020empirical), yet still struggle with long-range patterns and with modeling complex non-local interactions across economic drivers [ullah2025forecasting]. Methodologically, many studies emphasize technical improvements without rigorous comparison to the random walk benchmark [abedin2025deep], and often report only statistical metrics. Transaction-cost-adjusted evaluations remain scarce, leaving open whether model forecasts yield economically meaningful gains [galeshchuk2016neural, ayitey2023forex]. Determining whether any approach can outperform the random walk both statistically and economically remains a core question in daily exchange rate forecasting.
@@
The emergence of Transformer-based architectures offers a promising alternative [fischer2024fx], but their direct application to financial time series faces three critical limitations. First, the standard pointwise self-attention mechanism, designed for discrete tokens, fails to capture the continuous, trend-like features of financial data [zhang2025forecasting, wu2021autoformer]. Second, standard Transformers lack mechanisms to differentiate influential market drivers from redundant noise among correlated external factors [wang2024timexer, mishra2024volatility]. Third, these models remain black boxes. While post-hoc techniques like SHAP provide retrospective explanations, they fail to offer the forward-looking, time-varying interpretability required for adaptive investment decisions [bybee2023narrative, shi2025econometrics].
@@
We evaluate EXFormer using daily data for three major currency pairs (EUR/USD, USD/JPY, GBP/USD) alongside 28 macroeconomic and financial covariates. Our validation strategy is threefold: (1) we assess statistical accuracy using the Mean Squared Forecast Error (MSFE) ratio relative to the random walk, confirmed by the Clark–West test [clark2007approximately]; (2) we evaluate Directional Accuracy (DA) to verify the prediction of market turning pointsblaskowitz2014testing; and (3) we conduct a realistic trading backtest to bridge the gap between statistical gains and economic value, explicitly accounting for transaction costs.
@@ 2.1 The econometric models and the high-frequency challenge
The academic pursuit of exchange rate predictability is historically rooted in structural macroeconomic [forbes2018shocks]. Grounded in economic theory, models such as the monetary model and its sticky-price variant posit that exchange rate movements are driven by fundamentals like money supply, output levels, and inflation differentials [cheung2019exchange]. The subsequent research extended these to incorporate interest rate differentials, as formalized in uncovered interest parity [candian2025imperfect], and portfolio balance effects [camanho2022global]. While these models offer theoretical narratives, their empirical performance has been downgrade at daily horizons [bulut2018google].
The work of meese1983empirical crystallized this failure into what is now famously known as the ”Meese–Rogoff puzzle”. Their out-of-sample forecasting exercise demonstrated that a random walk model consistently outperformed a suite of structural models for major currencies at horizons of up to one year. A comprehensive review by cheung2005empirical re-affirmed the puzzle, concluding that no single model consistently and significantly dominates the random walk. The decades of following research have sought to resolve the puzzle. For example, mark2001nominal found evidence of long-horizon predictability using monetary fundamentals, but this finding was later challenged for its reliance on in-sample fitting and specific time periods [hong2024twofold]. Other research has explored nonlinearities, suggesting that transaction costs or policy interventions create thresholds beyond which fundamentals reassert their influence [rossi2013exchange, engel2019uncovered, flores2022forecasting, candian2025imperfect].
The macroeconomic variables central to these models are typically measured at low frequencies (weekly or monthly) [lilley2022exchange], making them inefficient to explain or predict the high-frequency, noisy dynamics of daily returns. Applying these models to the daily returns data often requires temporal disaggregation or interpolation techniques that introduce measurement error and model instabilitye [cheung2019exchange]. This leaves a research gap: the need for a framework that can integrate information across different time scales—from high-frequency market noise to slower-moving fundamental trends—without being constrained by the low-frequency nature of traditional economic data. Our EXFormer model addresses this directly through its multi-scale trend-aware attention mechanism, which is designed to analyze temporal patterns at various frequencies within a single, unified architecture.
@@ 2.2 The rise of machine learning and hybrid models
The limitations of linear econometric models prompted a shift towards more flexible, data-driven approaches from machine learning [mallqui2019predicting]. The early applications included SVMs and RFs, which demonstrated some success in capturing nonlinear relationships between predictors and exchange rate returns [fu2019evolutionary, gyamerah2020long]. However, they have treated financial data as a collection of independent and identically distributed feature vectors, largely ignore the temporal ordering of observations, failing to model path dependency, volatility clustering, and other sequential effects that are symbols of exchange rate time series [abedin2025deep]. Furthermore, the RNNs and their more advanced variants, LSTM and GRU networks incorporating feedback loops and memory cells, designed to capture sequential dependencies [fischer2018deep]. A growing body of literature has applied these architectures to exchange rates returns forecasting, with some studies reporting improved forecast accuracy over both linear models and machine learning methods [zhou2025forecasting, cao2020deep]. The study by [gu2020empirical] also demonstrated the power of deep learning in the broader context of asset pricing, showing they could identify risk premia more effectively than traditional methods.
@@
The classical deep learning architectures face persistent challenges when it comes to exchange rate returns forecasting [wei2018hybrid]. First, while LSTMs mitigate the vanishing gradient problem of simple RNNs, they can still struggle to capture very long-range dependencies, which are relevant for modeling multi-month cycles in currency markets [yildirim2021forecasting, chi2025enhancing]. Second, their strictly sequential processing mechanism is a bottleneck [mao2024unveiling]. They process information one time step at a time, making it difficult to model complex, non-adjacent interactions [han2024lest]. For example, a monetary policy announcement from 60 days prior might have a more direct impact on today’s return than a minor commodity price fluctuation from yesterday [shi2025econometrics]. Third, much of the existing literature suffers from a statistical versus economic significance gap [windsor2022improving, abedin2025deep]. Many studies report improvements in statistical error metrics without conducting rigorous backtesting [liu2024new]. As pointed out by rossi2013exchange, statistical outperformance does not guarantee economic value [darvas2025forecasting]. EXFormer is designed to overcome these specific limitations. Its Transformer-based core replaces sequential processing with a self-attention mechanism, enabling the model to directly assess relationships across all points in the input sequence simultaneously. Furthermore, we bridge the economic significance gap by conducting a trading backtests to validate the practical value of our forecasts, a step missing from much of the prior deep learning literature in this domain.
@@
Meanwhile, some of the literature concerns how exchange rate forecasts are evaluated. Much of the empirical studies of applying machine learning and deep learning have assessed performance through point forecasting error measures such as RMSE, MAE, and MSFE ratios [rossi2013exchange, darvas2025forecasting]. While informative about average accuracy, these symmetric, scale-dependent metrics are skeptical of sign, downside risk, and payoff asymmetries that matter for trading and risk management. Directional accuracy and associated inference have been used less systematically, and even when reported, they are not always linked to trading simulations or to risk-adjusted performance measures (e.g., Sharpe, Sortino, drawdowns), leaving a gap between statistical and economic significance. Recent work in artificial intelligence-driven finance has argued for objective functions that better reflect investment goals by penalizing costly errors more heavily. For example, the magnitude-aware directional losses [michankow2024mean, michankow2025alternative] weight sign mistakes by their return amplitudes, aligning optimization with realized trading outcomes rather than with squared-error fit alone. Related proposals incorporate asymmetry, tail sensitivity, or downside aversion into training objectives and evaluation metrics, improving the congruence between model learning and portfolio utility [windsor2022improving, abedin2025deep]. A further limitation of prior studies is the frequent absence of realistic backtesting (transaction costs, slippage, and liquidity), rolling re-estimation, and regime-aware analysis. When these ingredients are missing, reported gains can be overstated and fragile out of sample [rossi2013exchange]. Based on this background, we position our contribution along both MSFE-based measures and DA with robust inference, and we translate forecasts into implementable long–short strategies evaluated with risk-adjusted statistics and trading frictions. The evaluation perspective complements our architectural advances and narrows the gap between predictive accuracy and actionable performance in daily exchange rate markets.
@@
Furthermore, a parallel strand of the literature employs hybrid models for exchange rate returns forecasting. On the one hand is decomposition-ensemble frameworks [hao2023bi, henriquez2019combined, sun2018clustering, sun2019new]. The exchange rate prices or returns are first decomposed into trend, cyclical and residual components via empirical mode decomposition or wavelet transforms, then each component is forecast with a specialized network and recombined [wei2025wavelet]. These methods can isolate scale-specific patterns, yet they amplify model complexity, require ad-hoc frequency splits [zhou2025forecasting]. Furthermore, decomposition-ensemble approaches do not inherently address the dynamic relevance of external covariates spreads, volatility indices, commodity prices across time [ullah2025forecasting].
@@
On the other hand, hybrid models that combine different econometric models, machine learning, and deep learning approaches have also emerged [neghab2025explaining, petrosino2025garch, khashei2021kalman]. These typically involve using one model to generate features or signals that are then fed into another model for final prediction [plakandaras2015forecasting]. The rationale is that econometric models, such as GARCH-type volatility processes, cointegration frameworks, or error-correction mechanisms, embed structural domain knowledge and well-established statistical properties, while deep networks provide flexibility to capture nonlinearities and long-range dependencies [petrosino2025garch, khashei2021kalman, neghab2025explaining]. For example, ARIMA-LSTM or GARCH-RNN hybrids have been shown to improve short-horizon volatility and return forecasts by jointly modeling conditional mean and variance dynamics [plakandaras2015forecasting, wei2018hybrid]. While the methods can leverage the strengths of both paradigms, it often results in a loss of interpretability and does not fully exploit the temporal structure of the data [shi2025econometrics]. Furthermore, these methods have proven valuable in high-frequency daily data environments, where volatility clustering, regime switches, and heavy tails challenge either paradigm in isolation. However, hybrid frameworks often face trade-offs in interpretability, computational efficiency, and robustness out of sample [shi2025econometrics, ullah2025forecasting]. Our proposed EXFormer shares the hybrid motivation but differs in design. Instead of ad-hoc model stacking or decomposition, it embeds econometric intuition (trend extraction, multi-scale dependence) directly within the attention mechanism and feature selection modules, offering an integrated architecture that remains both scalable and interpretable for daily exchange rate forecasting.
@@ 2.3 Transformer models frontier
The Transformer architecture, first introduced by vaswani2017attention for natural language processing, has recently emerged as a powerful alternative for time series forecasting. Its core innovation, the self-attention mechanism, allows the model to weigh the influence of all other time steps when processing a given point in a sequence, addressing the long-range dependency and sequential bottleneck issues. Early applications in finance have shown promise [liu2024interpretable, mishra2024volatility], and have been developed to improve efficiency for very long sequence forecasting [wang2024timexer, wu2021autoformer]. However, traditional self-attention treats each time stamp as an exchangeable token and computes pointwise query-key similarity [zhang2025forecasting]. For noisy financial series, where whether an observation is an outlier, a break, or part of a pattern depends on its local trend, pointwise matching can misalign relevant contexts [hong2024twofold, han2024lest]. Recent decomposition or frequency domain Transformers partially mitigate this by separating trend, seasonality or filtering in the spectral domain, yet they do not explicitly encode local slope information into the attention kernel [fischer2024fx]. Moreover, exchange rate returns are influenced by a broad, time-varying set of exogenous drivers [stavrakeva2024fundamental, candian2025imperfect, iyke2022exchange]. The standard multi-head attention distributes capacity across heads but does not by itself distinguish influential covariates from redundant ones [mao2024unveiling]. The variable selection layers as in petrosino2025garch are a step forward, yet most implementations provide post-hoc importance or static gating, offering limited transparency about which inputs are used at prediction time. EXFormer is purpose-built to fill the gaps. The study aims to advance the state-of-the-art in the daily exchange rate returns forecasting.

### v2

@@ 1 Introduction
Exchange rates lie at the core of global finance, shaping trade, capital flows, monetary policy transmission, portfolio allocation, and corporate risk management [23]. Accurate day-ahead forecasts support hedging by multinational firms, guide central bank interventions, and inform systematic trading strategies used by hedge funds [62]. Yet despite decades of research, forecasting daily exchange rate returns remains one of the most persistent challenges in international finance [24, 63]. Meese and Rogoff [52] demonstrated that structural and time-series models routinely fail to outperform a simple random walk in out-of-sample tests. The Meese–Rogoff puzzle reflects two fundamental difficulties. First, daily returns combine high-frequency noise—microstructure effects, news-driven volatility, liquidity shocks—with slower patterns such as medium-term cycles from interest differentials and carry-trade activity [37, 56, 57, 50]. Second, it is necessary to model the influence and cross-market interactions of diverse exogenous drivers, such as other currency pairs, stock and commodity indices, and scheduled macroeconomic announcements on the daily returns behavior [68]. Traditional econometric models, especially those designed for lower-frequency data, often produce unstable daily forecasts and are prone to overfitting [18, 40]. Portfolio-balance theory further highlights the role of cross-asset capital flows and time-varying risk preferences, which conventional models typically overlook [20]. The core research gap is thus how to jointly model multi-scale temporal dependencies and multi-factor cross-market linkages. Although uncovered interest parity (UIP) implies unpredictability of exchange rate changes, its persistent empirical rejection indicates the presence of time-varying risk premia that more advanced forecasting models may exploit [23, 4].
@@
Recent work increasingly applies machine learning and deep learning to forecasting daily exchange rate returns [61, 71, 58]. Classical methods such as SVMs and random forests offer limited gains, as they ignore temporal structure [64, 34]. RNN-based models including LSTM and GRU networks, better capture sequential dependencies [82, 13] and have been widely tested in financial forecasting (Fischer and Krauss [25]; Gu et al. [32]), yet still struggle with long-range patterns and with modeling complex non-local interactions across economic drivers [72]. Methodologically, many studies emphasize technical improvements without rigorous comparison to the random walk benchmark [1], and often report only statistical metrics. Transaction-cost-adjusted evaluations remain scarce, leaving open whether model forecasts yield economically meaningful gains [31, 4]. Determining whether any approach can outperform the random walk both statistically and economically remains a core question in daily exchange rate forecasting.
@@
The emergence of Transformer-based architectures offers a promising alternative [26], but their direct application to financial time series faces three critical limitations. First, the standard pointwise self-attention mechanism, designed for discrete tokens, fails to capture the continuous, trend-like features of financial data [81, 79]. Second, standard Transformers lack mechanisms to differentiate influential market drivers from redundant noise among correlated external factors [75, 55]. Third, these models remain black boxes. While post-hoc techniques like SHAP provide retrospective explanations, they fail to offer the forward-looking, time-varying interpretability required for adaptive investment decisions [9, 65].
@@
We evaluate EXFormer using daily data for three major currency pairs (EUR/USD, USD/JPY, GBP/USD) alongside 28 macroeconomic and financial covariates. Our validation strategy is threefold: (1) we assess statistical accuracy using the Mean Squared Forecast Error (MSFE) ratio relative to the random walk, confirmed by the Clark–West test [17]; (2) we evaluate Directional Accuracy (DA) to verify the prediction of market turning pointsBlaskowitz and Herwartz [7]; and (3) we conduct a realistic trading backtest to bridge the gap between statistical gains and economic value, explicitly accounting for transaction costs.
@@ 2.1 The econometric models and the high-frequency challenge
The academic pursuit of exchange rate predictability is historically rooted in structural macroeconomic [28]. Grounded in economic theory, models such as the monetary model and its sticky-price variant posit that exchange rate movements are driven by fundamentals like money supply, output levels, and inflation differentials [14]. The subsequent research extended these to incorporate interest rate differentials, as formalized in uncovered interest parity [12], and portfolio balance effects [11]. While these models offer theoretical narratives, their empirical performance has been downgrade at daily horizons [8].
The work of Meese and Rogoff [52] crystallized this failure into what is now famously known as the ”Meese–Rogoff puzzle”. Their out-of-sample forecasting exercise demonstrated that a random walk model consistently outperformed a suite of structural models for major currencies at horizons of up to one year. A comprehensive review by Cheung et al. [15] re-affirmed the puzzle, concluding that no single model consistently and significantly dominates the random walk. The decades of following research have sought to resolve the puzzle. For example, Mark and Sul [51] found evidence of long-horizon predictability using monetary fundamentals, but this finding was later challenged for its reliance on in-sample fitting and specific time periods [39]. Other research has explored nonlinearities, suggesting that transaction costs or policy interventions create thresholds beyond which fundamentals reassert their influence [63, 20, 27, 12].
The macroeconomic variables central to these models are typically measured at low frequencies (weekly or monthly) [46], making them inefficient to explain or predict the high-frequency, noisy dynamics of daily returns. Applying these models to the daily returns data often requires temporal disaggregation or interpolation techniques that introduce measurement error and model instabilitye [14]. This leaves a research gap: the need for a framework that can integrate information across different time scales—from high-frequency market noise to slower-moving fundamental trends—without being constrained by the low-frequency nature of traditional economic data. Our EXFormer model addresses this directly through its multi-scale trend-aware attention mechanism, which is designed to analyze temporal patterns at various frequencies within a single, unified architecture.
@@ 2.2 The rise of machine learning and hybrid models
The limitations of linear econometric models prompted a shift towards more flexible, data-driven approaches from machine learning [49]. The early applications included SVMs and RFs, which demonstrated some success in capturing nonlinear relationships between predictors and exchange rate returns [29, 34]. However, they have treated financial data as a collection of independent and identically distributed feature vectors, largely ignore the temporal ordering of observations, failing to model path dependency, volatility clustering, and other sequential effects that are symbols of exchange rate time series [1]. Furthermore, the RNNs and their more advanced variants, LSTM and GRU networks incorporating feedback loops and memory cells, designed to capture sequential dependencies [25]. A growing body of literature has applied these architectures to exchange rates returns forecasting, with some studies reporting improved forecast accuracy over both linear models and machine learning methods [82, 13]. The study by [32] also demonstrated the power of deep learning in the broader context of asset pricing, showing they could identify risk premia more effectively than traditional methods.
@@
The classical deep learning architectures face persistent challenges when it comes to exchange rate returns forecasting [77]. First, while LSTMs mitigate the vanishing gradient problem of simple RNNs, they can still struggle to capture very long-range dependencies, which are relevant for modeling multi-month cycles in currency markets [80, 16]. Second, their strictly sequential processing mechanism is a bottleneck [50]. They process information one time step at a time, making it difficult to model complex, non-adjacent interactions [35]. For example, a monetary policy announcement from 60 days prior might have a more direct impact on today’s return than a minor commodity price fluctuation from yesterday [65]. Third, much of the existing literature suffers from a statistical versus economic significance gap [78, 1]. Many studies report improvements in statistical error metrics without conducting rigorous backtesting [48]. As pointed out by Rossi [63], statistical outperformance does not guarantee economic value [18]. EXFormer is designed to overcome these specific limitations. Its Transformer-based core replaces sequential processing with a self-attention mechanism, enabling the model to directly assess relationships across all points in the input sequence simultaneously. Furthermore, we bridge the economic significance gap by conducting a trading backtests to validate the practical value of our forecasts, a step missing from much of the prior deep learning literature in this domain.
@@
Meanwhile, some of the literature concerns how exchange rate forecasts are evaluated. Much of the empirical studies of applying machine learning and deep learning have assessed performance through point forecasting error measures such as RMSE, MAE, and MSFE ratios [63, 18]. While informative about average accuracy, these symmetric, scale-dependent metrics are skeptical of sign, downside risk, and payoff asymmetries that matter for trading and risk management. Directional accuracy and associated inference have been used less systematically, and even when reported, they are not always linked to trading simulations or to risk-adjusted performance measures (e.g., Sharpe, Sortino, drawdowns), leaving a gap between statistical and economic significance. Recent work in artificial intelligence-driven finance has argued for objective functions that better reflect investment goals by penalizing costly errors more heavily. For example, the magnitude-aware directional losses [53, 54] weight sign mistakes by their return amplitudes, aligning optimization with realized trading outcomes rather than with squared-error fit alone. Related proposals incorporate asymmetry, tail sensitivity, or downside aversion into training objectives and evaluation metrics, improving the congruence between model learning and portfolio utility [78, 1]. A further limitation of prior studies is the frequent absence of realistic backtesting (transaction costs, slippage, and liquidity), rolling re-estimation, and regime-aware analysis. When these ingredients are missing, reported gains can be overstated and fragile out of sample [63]. Based on this background, we position our contribution along both MSFE-based measures and DA with robust inference, and we translate forecasts into implementable long–short strategies evaluated with risk-adjusted statistics and trading frictions. The evaluation perspective complements our architectural advances and narrows the gap between predictive accuracy and actionable performance in daily exchange rate markets.
@@
Furthermore, a parallel strand of the literature employs hybrid models for exchange rate returns forecasting. On the one hand is decomposition-ensemble frameworks [36, 38, 69, 70]. The exchange rate prices or returns are first decomposed into trend, cyclical and residual components via empirical mode decomposition or wavelet transforms, then each component is forecast with a specialized network and recombined [76]. These methods can isolate scale-specific patterns, yet they amplify model complexity, require ad-hoc frequency splits [82]. Furthermore, decomposition-ensemble approaches do not inherently address the dynamic relevance of external covariates spreads, volatility indices, commodity prices across time [72].
@@
On the other hand, hybrid models that combine different econometric models, machine learning, and deep learning approaches have also emerged [58, 60, 42]. These typically involve using one model to generate features or signals that are then fed into another model for final prediction [61]. The rationale is that econometric models, such as GARCH-type volatility processes, cointegration frameworks, or error-correction mechanisms, embed structural domain knowledge and well-established statistical properties, while deep networks provide flexibility to capture nonlinearities and long-range dependencies [60, 42, 58]. For example, ARIMA-LSTM or GARCH-RNN hybrids have been shown to improve short-horizon volatility and return forecasts by jointly modeling conditional mean and variance dynamics [61, 77]. While the methods can leverage the strengths of both paradigms, it often results in a loss of interpretability and does not fully exploit the temporal structure of the data [65]. Furthermore, these methods have proven valuable in high-frequency daily data environments, where volatility clustering, regime switches, and heavy tails challenge either paradigm in isolation. However, hybrid frameworks often face trade-offs in interpretability, computational efficiency, and robustness out of sample [65, 72]. Our proposed EXFormer shares the hybrid motivation but differs in design. Instead of ad-hoc model stacking or decomposition, it embeds econometric intuition (trend extraction, multi-scale dependence) directly within the attention mechanism and feature selection modules, offering an integrated architecture that remains both scalable and interpretable for daily exchange rate forecasting.
@@ 2.3 Transformer models frontier
The Transformer architecture, first introduced by Vaswani et al. [73] for natural language processing, has recently emerged as a powerful alternative for time series forecasting. Its core innovation, the self-attention mechanism, allows the model to weigh the influence of all other time steps when processing a given point in a sequence, addressing the long-range dependency and sequential bottleneck issues. Early applications in finance have shown promise [47, 55], and have been developed to improve efficiency for very long sequence forecasting [75, 79]. However, traditional self-attention treats each time stamp as an exchangeable token and computes pointwise query-key similarity [81]. For noisy financial series, where whether an observation is an outlier, a break, or part of a pattern depends on its local trend, pointwise matching can misalign relevant contexts [39, 35]. Recent decomposition or frequency domain Transformers partially mitigate this by separating trend, seasonality or filtering in the spectral domain, yet they do not explicitly encode local slope information into the attention kernel [26]. Moreover, exchange rate returns are influenced by a broad, time-varying set of exogenous drivers [68, 12, 40]. The standard multi-head attention distributes capacity across heads but does not by itself distinguish influential covariates from redundant ones [50]. The variable selection layers as in Petrosino et al. [60] are a step forward, yet most implementations provide post-hoc importance or static gating, offering limited transparency about which inputs are used at prediction time. EXFormer is purpose-built to fill the gaps. The study aims to advance the state-of-the-art in the daily exchange rate returns forecasting.